# Branch structure and nonextensive thermodynamics of Kalb-Ramond-ModMax black holes: observational signatures

### v1

@@ Abstract
We investigate a static, spherically symmetric black hole arising in Einstein gravity coupled to a Kalb-Ramond field and ModMax nonlinear electrodynamics, both of which are independently well motivated extensions of standard electrovacuum gravity. The solution depends, beyond mass and charge, on a Lorentz-violating parameter, a ModMax deformation parameter, and a discrete branch selector $\zeta=\pm 1$. We show that the ordinary branch admits extremal and non-extremal configurations, while the phantom branch generically supports a single-horizon geometry. Black-hole thermodynamics is analyzed within the Tsallis non-extensive framework, revealing branch-dependent stability and Joule-Thomson behavior. Weak gravitational lensing, photon propagation in plasma, and tidal forces are then studied, revealing clear optical and strong-field signatures that distinguish the two branches. In particular, the ordinary branch exhibits finite-radius tidal inversion, absent in the phantom sector. Our results demonstrate that the combined Kalb-Ramond and ModMax effects lead to a rich and observationally distinguishable black-hole phenomenology.
@@ I Introduction
On the observational side, gravitational lensing remains one of the most direct probes of spacetime geometry in the strong-field regime. Since its first experimental confirmation Dyson et al. (1920), lensing has evolved into a precision tool for testing GR and constraining alternative theories Will (2014). The Gauss-Bonnet (GB) approach Gibbons and Werner (2008); Werner (2012) offers a coordinate-independent method for computing weak-field deflection angles and has been widely applied in BH spacetimes Ren et al. (2021); Sucu and Övgün (2024); Sucu (2026); Sucu and Sakalli (2025b); Wang et al. (2024). Moreover, realistic astrophysical environments are typically permeated by plasma, which modifies photon propagation through dispersive effects. In such settings, the photon sphere and shadow properties depend not only on the spacetime geometry but also on the plasma distribution, providing additional discriminants among competing gravitational models Perlick et al. (2015); Rogers (2015); Sucu et al. (2025).
@@ V Gravitational lensing and photon spheres
In the present context, weak gravitational lensing offers a sensitive diagnostic of the Kalb-Ramond-ModMax black hole. Although both the Kalb-Ramond condensate and the ModMax nonlinearity primarily affect the near-horizon geometry, their influence propagates to large distances through the optical structure of spacetime. To capture these effects in a coordinate-independent manner, we employ the Gauss-Bonnet (GB) theorem approach Gibbons and Werner (2008); Sucu and Sakalli (2025c); Sucu and Sakallı (2025c), which interprets the deflection of light as a global geometric effect governed by the integrated Gaussian curvature of an associated optical manifold.
V.1 Optical geometry and Gauss-Bonnet method
In static and spherically symmetric spacetimes, null geodesics can be equivalently described as spatial geodesics on an effective two-dimensional Riemannian manifold known as the optical geometry ABRAMOWICZ et al. (1988); Perlick (2006). Restricting the Kalb-Ramond-ModMax metric (1) to the equatorial plane ($\theta=\pi/2$), the corresponding optical line element takes the form
@@ V.1 Optical geometry and Gauss-Bonnet method
The intrinsic curvature of the optical manifold is characterized by the Gaussian curvature $K$, which in two dimensions is fully determined by a single independent component of the Riemann tensor Sucu and Sakallı (2025a), namely
$K=\frac{R_{r\phi r\phi}}{\det(g_{\rm opt})}.$ (26)
In the weak-field region ($r\gg r_{h}$), we expand the metric function asymptotically and retain terms up to $\mathcal{O}(r^{-5})$. This yields the following expression for the Gaussian curvature:
$\!\!\!\!\!K\simeq\frac{3\zeta Q^{2}}{r^{4}(1-\ell)^{4}e^{\gamma}}-\frac{3\zeta Q^{2}\ell}{r^{4}(1-\ell)^{4}e^{\gamma}}+\frac{2M\ell^{3}}{r^{3}(1-\ell)^{4}}$
$\ \ \ \ -\frac{6M\ell^{2}}{r^{3}(1-\ell)^{4}}+\frac{6M\ell}{r^{3}(1-\ell)^{4}}-\frac{2M}{r^{3}(1-\ell)^{4}}$
$\ \ \ \ +\frac{3M^{2}\ell^{4}}{r^{4}(1-\ell)^{4}}-\frac{12M^{2}\ell^{3}}{r^{4}(1-\ell)^{4}}+\frac{18M^{2}\ell^{2}}{r^{4}(1-\ell)^{4}}$
$\ \ \ \ -\frac{12M^{2}\ell}{r^{4}(1-\ell)^{4}}+\frac{3M^{2}}{r^{4}(1-\ell)^{4}}.$ (27)
The leading $r^{-3}$ contributions originate from the mass sector, while the $r^{-4}$ terms encode both second-order mass corrections and the charge-dependent ModMax contributions, controlled by the branch parameter $\zeta$ and the suppression factor $e^{-\gamma}$. The Kalb-Ramond parameter $\ell$ enters universally through the prefactor $(1-\ell)^{-4}$, amplifying all curvature contributions as $\ell$ approaches unity.
@@
The Gauss-Bonnet theorem provides a direct link between this intrinsic curvature and the total deflection angle experienced by a light ray. Applied to a non-singular domain $\mathcal{D}_{\infty}$ exterior to the photon trajectory, the theorem yields the coordinate-independent expression Gibbons and Werner (2008)
$\beta=-\iint_{\mathcal{D}_{\infty}}K\,dS,$ (28)
where $dS$ denotes the surface element of the optical manifold. In the weak-lensing approximation, the photon path is well approximated by the unperturbed straight-line trajectory $r=b/\sin\phi$, allowing the integral to be evaluated analytically. The resulting deflection angle, together with its branch-dependent structure and limiting behavior, is discussed in the following subsection.
V.2 Weak-field deflection angle and branch-dependent signatures
Within the Gauss-Bonnet framework, the total deflection angle experienced by a light ray with impact parameter $b$ is determined by the surface integral of the Gaussian curvature over the exterior optical domain $\mathcal{D}_{\infty}$, namely
$\beta=-\iint_{\mathcal{D}_{\infty}}K\,dS,$ (29)
where the integration region is bounded by the photon trajectory and an asymptotic reference curve Gibbons and Werner (2008). In the weak-field regime, the photon path is well approximated by the unperturbed straight-line trajectory $r=b/\sin\phi$, which allows the integral to be evaluated analytically.
@@ V.2 Weak-field deflection angle and branch-dependent signatures
Carrying out the integration using the asymptotic expansion of the Gaussian curvature given in Eq. (27), we obtain the vacuum deflection angle for the Kalb-Ramond-ModMax black hole,
$\!\!\!\!\!\!\beta_{\rm vac}\approx\frac{4M}{b(1-\ell)^{4}}-\frac{12M\ell}{b(1-\ell)^{4}}+\frac{12M\ell^{2}}{b(1-\ell)^{4}}-\frac{4M\ell^{3}}{b(1-\ell)^{4}}$
$\!\!\!\!\!\!+\frac{15\pi M^{2}}{4b^{2}(1-\ell)^{4}}-\frac{3\pi M^{2}\ell}{2b^{2}(1-\ell)^{4}}+\frac{3\pi M^{2}\ell^{3}}{2b^{2}(1-\ell)^{4}}-\frac{3\pi M^{2}\ell^{4}}{4b^{2}(1-\ell)^{4}}$
$\!\!\!\!\!\!-\frac{3\pi Q^{2}\zeta}{4b^{2}(1-\ell)^{4}e^{\gamma}}+\frac{3\pi Q^{2}\zeta\ell}{4b^{2}(1-\ell)^{4}e^{\gamma}}-\frac{4MQ^{2}\zeta}{3b^{3}(1-\ell)^{4}e^{\gamma}}$
$\!\!\!\!\!\!+\frac{8MQ^{2}\ell^{2}\zeta}{3b^{3}(1-\ell)^{4}e^{\gamma}}-\frac{4MQ^{2}\ell\zeta}{3b^{3}(1-\ell)^{4}e^{\gamma}}+\mathcal{O}(M^{3},1/b^{4}).$ (30)
This expression naturally decomposes into three classes of contributions: (i) mass terms proportional to $M/b$ and $M^{2}/b^{2}$, (ii) Kalb-Ramond corrections controlled by powers of $\ell$ and the universal amplification factor $(1-\ell)^{-4}$, and (iii) charge-dependent ModMax terms proportional to $Q^{2}\zeta$, whose magnitude is suppressed by $e^{-\gamma}$ and whose sign is fixed by the branch parameter $\zeta$.
@@
$\beta_{\rm Sch}=\frac{4M}{b}+\frac{15\pi M^{2}}{4b^{2}}+\mathcal{O}(b^{-3}),$ (31)
recovering the standard general-relativistic result Epstein and Shapiro (1980); Weinberg (2013). For the ordinary branch $(\zeta=+1)$ with $\ell\to 0$ and $\gamma\to 0$, the expression reduces to the Reissner-Nordström deflection angle, i.e.
$\beta_{\rm RN}=\frac{4M}{b}+\frac{15\pi M^{2}}{4b^{2}}-\frac{3\pi Q^{2}}{4b^{2}}-\frac{4MQ^{2}}{3b^{3}}+\mathcal{O}(b^{-4}),$ (32)
@@
On the other hand, in the phantom branch $(\zeta=-1)$ the sign reversal of the charge sector transforms the electromagnetic contribution into an additional focusing mechanism. Even in the limit $\ell\to 0$ and $\gamma\to 0$, the deflection angle becomes
$\beta_{\rm phantom}=\frac{4M}{b}+\frac{15\pi M^{2}}{4b^{2}}+\frac{3\pi Q^{2}}{4b^{2}}+\frac{4MQ^{2}}{3b^{3}}+\mathcal{O}(b^{-4}),$ (33)
which has no analogue in standard general relativity. In this sector, both mass and charge contribute constructively to light bending, leading to systematically larger deflection angles.
@@
Figure 8: Vacuum deflection angle $\beta_{\rm vac}$ as a function of the impact parameter $b$ for KR-ModMax black holes with fixed $M=1$, $Q=0.5$, and $\gamma=2$. The color scale indicates the Kalb-Ramond parameter $\ell\in[0,0.5]$. Panel (a) shows the ordinary branch ($\zeta=+1$), where the electromagnetic contribution partially counteracts gravitational focusing, leading to suppressed deflection at small $b$. Increasing $\ell$ enhances the mass-dominated terms through the $(1-\ell)^{-4}$ amplification. Panel (b) displays the phantom branch ($\zeta=-1$), where the sign reversal of the electromagnetic sector transforms the charge contribution into an additional focusing mechanism. Consequently, the deflection angle remains strictly positive and is notably enhanced at small $b$, exceeding the ordinary-branch values by up to an order of magnitude as $\ell$ increases. This qualitative distinction serves as an observational discriminator between the two electromagnetic sectors.
These qualitative differences are evident in the numerical results shown in Fig. 8, where the deflection angle is plotted for varying Kalb-Ramond parameter $\ell\in[0,0.5]$ at fixed $\gamma=2$. In the ordinary branch, the competition between mass-induced focusing and charge-induced defocusing produces smaller deflection angles, particularly at low $\ell$. As $\ell$ increases, the universal amplification factor $(1-\ell)^{-4}$ enhances all contributions, leading to progressively larger deflection at all impact parameters. In the phantom branch, the deflection angle remains strictly positive and is substantially enhanced relative to the ordinary branch, exceeding the latter by more than an order of magnitude at small $b$ for larger $\ell$ values. This pronounced enhancement provides a clear geometric manifestation of the branch dichotomy.
From an observational perspective, these branch-dependent signatures translate into potentially measurable effects. Phantom configurations would produce significantly larger Einstein ring radii and enhanced magnification in strong-lensing systems, while the $\ell$-dependence of the weak-field deflection could imprint small but systematic astrometric shifts in high-precision measurements of stellar orbits near supermassive black holes. Together with the thermodynamic and photon-sphere properties discussed in previous sections, gravitational lensing offers a complementary channel for discriminating between the ordinary and phantom Kalb-Ramond-ModMax black hole branches.

### v2

@@ Abstract
We investigate a static, spherically symmetric black hole arising in Einstein gravity coupled to a Kalb-Ramond field and ModMax nonlinear electrodynamics, both of which are independently well motivated extensions of standard electrovacuum gravity. The solution depends, beyond mass and charge, on a Lorentz-violating parameter, a ModMax deformation parameter, and a discrete branch selector $\zeta=\pm 1$. We show that the ordinary branch admits extremal and non-extremal configurations, while the phantom branch generically supports a single-horizon geometry. Black-hole thermodynamics is analyzed within the Tsallis non-extensive framework, revealing branch-dependent stability and Joule-Thomson behavior. Weak gravitational lensing is analyzed using the Ono-Ishihara-Asada extension of the Gauss-Bonnet theorem, which is required by the non-Euclidean asymptotic structure of the Kalb-Ramond optical geometry and yields a negative topological correction that reduces light bending relative to the Schwarzschild baseline. Photon propagation in plasma and tidal forces are also studied, revealing clear optical and strong-field signatures that distinguish the two branches.
@@ I Introduction
On the observational side, gravitational lensing remains one of the most direct probes of spacetime geometry in the strong-field regime. Since its first experimental confirmation Dyson et al. (1920), lensing has evolved into a precision tool for testing GR and constraining alternative theories Will (2014). The Gauss-Bonnet (GB) approach Gibbons and Werner (2008); Werner (2012) offers a coordinate-independent method for computing weak-field deflection angles and has been widely applied in BH spacetimes Ren et al. (2021); Sucu and Övgün (2024); Sucu (2026); Sucu and Sakalli (2025b); Wang et al. (2024). However, the standard formulation assumes asymptotically Euclidean optical geometry, a condition that is violated for the KR-ModMax metric due to the modified lapse at infinity. In such cases, the Ono-Ishihara-Asada (OIA) generalisation Ono et al. (2017); Ishihara et al. (2016); Jusufi et al. (2017) must be employed, which introduces a topological boundary correction encoding the non-flat asymptotic structure. Moreover, realistic astrophysical environments are typically permeated by plasma, which modifies photon propagation through dispersive effects. In such settings, the photon sphere and shadow properties depend not only on the spacetime geometry but also on the plasma distribution, providing additional discriminants among competing gravitational models Perlick et al. (2015); Rogers (2015); Sucu et al. (2025).
@@ V Gravitational lensing and photon spheres
In the present context, weak gravitational lensing offers a sensitive diagnostic of the Kalb-Ramond-ModMax black hole. Although both the Kalb-Ramond condensate and the ModMax nonlinearity primarily affect the near-horizon geometry, their influence propagates to large distances through the optical structure of spacetime. A key feature of the metric (3) is its non-standard asymptotic behaviour, $f(r)\to 1/(1-\ell)\neq 1$ as $r\to\infty$, which renders the standard Gibbons-Werner formulation of the Gauss-Bonnet theorem inapplicable Gibbons and Werner (2008); Jusufi et al. (2017); Ishihara et al. (2016). Instead, the deflection angle must be computed using the Ono-Ishihara-Asada (OIA) extension Ono et al. (2017); Ishihara et al. (2016), which incorporates a topological boundary correction arising from the non-Euclidean asymptotic structure of the optical manifold.
V.1 Non-asymptotic flatness and coordinate transformation
The optical metric associated with a static, spherically symmetric spacetime (1) restricted to the equatorial plane ($\theta=\pi/2$) takes the form
@@ V.1 Non-asymptotic flatness and coordinate transformation
As noted in Eq. (4), the metric function satisfies $f(r)\to 1/(1-\ell)$ as $r\to\infty$. As a consequence, the geodesic curvature of the circular boundary $C_{R}$ at large $R$ in the optical geometry evaluates to
$\lim_{R\to\infty}\kappa(C_{R})\,\frac{dt}{d\phi}=\sqrt{f_{\infty}}=\frac{1}{\sqrt{1-\ell}}\neq 1,$ (26)
confirming that the optical manifold is not asymptotically Euclidean Jusufi et al. (2017); Ono et al. (2017). Therefore, the standard Gibbons-Werner Gauss-Bonnet formula Gibbons and Werner (2008) must be replaced by the OIA generalisation.
To make the structure transparent, we define $K\equiv\sqrt{1-\ell}$ and introduce the rescaled metric function
$h(r)\;\equiv\;K^{2}\,f(r)=1-\frac{2M(1-\ell)}{r}+\frac{\zeta\,Q^{2}\,e^{-\gamma}}{(1-\ell)\,r^{2}},$ (27)
which by construction satisfies $h(r)\to 1$ as $r\to\infty$. Under the coordinate transformation $\tilde{t}=t/K$, $\tilde{r}=K\,r$, the line element (1) takes the Barriola-Vilenkin-type form Jusufi et al. (2017)
$ds^{2}=-\hat{h}(\tilde{r})\,d\tilde{t}^{2}+\frac{d\tilde{r}^{2}}{\hat{h}(\tilde{r})}+\frac{\tilde{r}^{2}}{K^{2}}\,d\Omega^{2},$ (28)
where
$\hat{h}(\tilde{r})=1-\frac{2\hat{M}}{\tilde{r}}+\frac{\hat{Q}^{2}}{\tilde{r}^{2}},$ (29)
with the effective parameters
$\hat{M}=M(1-\ell)^{3/2},\qquad\hat{Q}^{2}=\zeta\,Q^{2}\,e^{-\gamma}.$ (30)
The angular part $\tilde{r}^{2}/K^{2}$ in Eq. (28) shows that the transformed metric possesses an angular surplus (since $1/K^{2}>1$ for $\ell>0$), in contrast to the angular deficit arising in the Letelier cloud-of-strings or global monopole spacetimes studied in Jusufi et al. (2017); Sucu and Sakalli (2025c). The physical impact parameter in the rescaled frame is $\tilde{b}=Kb$, where $b$ is the coordinate distance of closest approach in the original frame. The rescaled function $\hat{h}$ has exactly the Reissner-Nordström form, so all standard weak-field lensing results can be imported once the OIA boundary correction is included.
V.2 Gauss-Bonnet method with OIA correction
@@
The Gauss-Bonnet theorem applied to the non-singular domain $\mathcal{D}_{\infty}$ exterior to the photon trajectory in the optical geometry yields Gibbons and Werner (2008); Jusufi et al. (2017); Ono et al. (2017)
$\iint_{\mathcal{D}_{\infty}}\!\mathcal{K}\,dS+\frac{1}{K}\bigl(\pi+\hat{\alpha}\bigr)=\pi,$ (31)
where $\mathcal{K}$ is the Gaussian curvature and $dS$ is the surface element of the optical manifold, and the factor $1/K$ arises from Eq. (26). Solving for the deflection angle gives
$\hat{\alpha}=\pi(K-1)-K\iint_{\mathcal{D}_{\infty}}\!\mathcal{K}\,dS.$ (32)
The first term, $\pi(\sqrt{1-\ell}-1)<0$, is the topological boundary correction absent in the standard Gibbons-Werner approach. For the Letelier or global monopole spacetimes, this correction is positive and proportional to the angular deficit Jusufi et al. (2017); in the present case the sign is reversed because the KR condensate modifies the lapse rather than the angular metric.
The Gaussian curvature of the optical metric (25) is determined by Sucu and Sakallı (2025a)
$\mathcal{K}=-\frac{f^{\prime\,2}-2f\,f^{\prime\prime}}{4},$ (33)
which in the weak-field region ($r\gg r_{h}$) yields
$\mathcal{K}\simeq-\frac{2M}{(1-\ell)\,r^{3}}+\frac{3M^{2}+\dfrac{3\zeta\,Q^{2}\,e^{-\gamma}}{(1-\ell)^{3}}}{r^{4}}+\mathcal{O}(r^{-5}).$ (34)
The leading $r^{-3}$ contribution originates from the mass sector and carries a factor $(1-\ell)^{-1}$ reflecting the renormalized gravitational coupling. The $r^{-4}$ terms encode both second-order mass corrections and the charge-dependent ModMax contributions; the latter are suppressed by $e^{-\gamma}$ and amplified by $(1-\ell)^{-3}$.
To evaluate the curvature integral in Eq. (32) to $\mathcal{O}(M^{2}/b^{2})$ accuracy, the straight-line trajectory $r=b/\sin\phi$ is insufficient for the $M^{2}$ coefficient Pereira and others (2025); Araújo Filho and others (2025). Instead, one employs the iterative photon trajectory
$\frac{1}{\tilde{r}}=\frac{\sin\phi}{\tilde{b}}+\frac{\hat{M}\,(1-\cos\phi)^{2}}{\tilde{b}^{2}},$ (35)
which accounts for the self-consistent bending of the photon path to first post-Minkowskian order Pereira and others (2025). With this improved trajectory, the curvature integral yields the correct coefficient $15\pi/(4b^{2})$ for the $M^{2}$ contribution, replacing the lower value $3\pi/(4b^{2})$ obtained from the straight-line approximation.
V.3 Weak-field deflection angle and branch-dependent signatures
@@
Carrying out the integration in Eq. (32) with the iterative trajectory (35) and the Gaussian curvature (34), we obtain the vacuum deflection angle for the Kalb-Ramond-ModMax black hole:
$\hat{\alpha}_{\rm vac}$ $=$ $\pi\!\left(\sqrt{1-\ell}\,-1\right)+\frac{4M(1-\ell)}{b}$ (36)
$\!\!\!\!\!\!+\frac{\pi\!\left[15M^{2}(1-\ell)^{2}-\dfrac{3\zeta\,Q^{2}\,e^{-\gamma}}{1-\ell}\right]}{4b^{2}}$
$\!\!\!\!\!\!-\frac{4M\zeta\,Q^{2}\,e^{-\gamma}}{3b^{3}}+\mathcal{O}(b^{-4}).$
This expression naturally decomposes into four physically distinct contributions: (i) the topological boundary term $\pi(\sqrt{1-\ell}-1)<0$, which arises from the non-Euclidean asymptotic structure and reduces the total deflection; (ii) the leading mass term $4M(1-\ell)/b$, which is suppressed relative to the Schwarzschild value $4M/b$; (iii) second-order terms at $\mathcal{O}(b^{-2})$ encoding gravitational self-interaction and charge-dependent ModMax corrections; and (iv) a mixed mass-charge term at $\mathcal{O}(b^{-3})$, which is independent of $\ell$ and coincides with the standard Reissner-Nordström expression.
Expanding to first order in $\ell$, which is the physically relevant regime given observational bounds on Lorentz-violating parameters Masood (2024), the deflection angle simplifies to
$\hat{\alpha}_{\rm vac}$ $\approx$ $-\frac{\pi\,\ell}{2}+\frac{4M}{b}\left(1-\ell\right)$ (37)
$+\frac{\pi\!\left(15M^{2}-3\zeta\,Q^{2}\,e^{-\gamma}\right)}{4b^{2}}+\mathcal{O}(\ell^{2},\ell/b^{2}),$
where the $\ell$-dependence enters exclusively through the boundary term and the leading mass contribution.
@@
$\hat{\alpha}_{\rm Sch}=\frac{4M}{b}+\frac{15\pi M^{2}}{4b^{2}}+\mathcal{O}(b^{-3}),$ (38)
recovering the standard general-relativistic result including the correct second-order coefficient Epstein and Shapiro (1980); Weinberg (2013); Pereira and others (2025). For the ordinary branch $(\zeta=+1)$ with $\ell\to 0$ and $\gamma\to 0$, the expression reduces to the Reissner-Nordström deflection angle, i.e.
$\hat{\alpha}_{\rm RN}=\frac{4M}{b}+\frac{15\pi M^{2}}{4b^{2}}-\frac{3\pi Q^{2}}{4b^{2}}-\frac{4MQ^{2}}{3b^{3}}+\mathcal{O}(b^{-4}),$ (39)
@@
In the phantom branch $(\zeta=-1)$ with $\ell\to 0$ and $\gamma\to 0$, the sign reversal of the charge sector transforms the electromagnetic contribution into an additional focusing mechanism:
$\hat{\alpha}_{\rm phantom}=\frac{4M}{b}+\frac{15\pi M^{2}}{4b^{2}}+\frac{3\pi Q^{2}}{4b^{2}}+\frac{4MQ^{2}}{3b^{3}}+\mathcal{O}(b^{-4}),$ (40)
where both mass and charge contribute constructively to light bending, producing systematically larger deflection angles.
@@
Figure 8: Vacuum deflection angle $\hat{\alpha}_{\rm vac}$ as a function of the impact parameter $b$ for KR-ModMax black holes with fixed $M=1$, $Q=0.5$, and $\gamma=2$. The colour scale indicates the Kalb-Ramond parameter $\ell\in[0,0.5]$. Panel (a) shows the ordinary branch ($\zeta=+1$), where the electromagnetic contribution partially counteracts gravitational focusing. Increasing $\ell$ reduces the total deflection through the negative boundary term $\pi(\sqrt{1-\ell}-1)$ and the suppressed effective mass $M_{\rm eff}=M(1-\ell)$. Panel (b) displays the phantom branch ($\zeta=-1$), where the sign reversal of the electromagnetic sector enhances focusing. Although the phantom-branch deflection angles remain larger than their ordinary counterparts, increasing $\ell$ reduces both branches relative to their Schwarzschild baselines through the universal topological correction.
The qualitative features of Eq. (36) are illustrated in Fig. 8. A distinctive prediction of the OIA-corrected analysis is that the Kalb-Ramond parameter $\ell$ reduces the deflection angle through two mechanisms: the negative topological boundary contribution $\pi(\sqrt{1-\ell}-1)$, and the suppression of the effective gravitational mass to $M(1-\ell)$ in the leading term. This contrasts qualitatively with the Letelier cloud-of-strings spacetime, where the analogous topological term is positive and enhances deflection Jusufi et al. (2017). The difference is rooted in the distinct locations of non-flatness: the Letelier string cloud modifies the angular metric ($g_{\phi\phi}=(1-A)r^{2}$), producing an angular deficit, while the KR condensate modifies the lapse ($f_{\infty}=1/(1-\ell)$), producing an effective angular surplus in the optical geometry.
The branch dichotomy persists: the phantom branch ($\zeta=-1$) yields larger deflection angles than the ordinary branch ($\zeta=+1$) at fixed $\ell$, due to the constructive alignment of mass and charge contributions. However, both branches experience the same $\ell$-dependent reduction, so the ratio of phantom-to-ordinary deflection angles remains sensitive to the charge-to-mass ratio and the ModMax parameter $\gamma$.
From an observational perspective, the reduction of the deflection angle with increasing $\ell$ would manifest as smaller Einstein ring radii and decreased magnification in strong-lensing systems compared to the Schwarzschild prediction at the same mass, providing a distinctive observational signature of Kalb-Ramond Lorentz symmetry breaking. Together with the thermodynamic and photon-sphere properties discussed in previous sections, gravitational lensing offers a complementary channel for constraining the KR parameter $\ell$ and discriminating between the ordinary and phantom Kalb-Ramond-ModMax black hole branches.